# Red-Teaming Vision-Language-Action Models via Quality Diversity Prompt Generation for Robust Robot Policies

### 12 Mar 2026 v1

@@ Abstract
Vision-Language-Action (VLA) models have significant potential to enable general-purpose robotic systems for a range of vision-language tasks. However, the performance of VLA-based robots is highly sensitive to the precise wording of language instructions, and it remains difficult to predict when such robots will fail. To improve the robustness of VLAs to different wordings, we present Q-DIG (Quality Diversity for Diverse Instruction Generation), which performs red-teaming by scalably identifying diverse natural language task descriptions that induce failures while remaining task-relevant. Q-DIG integrates Quality Diversity (QD) techniques with Vision-Language Models (VLMs) to generate a broad spectrum of adversarial instructions that expose meaningful vulnerabilities in VLA behavior. Our results across multiple simulation benchmarks show that Q-DIG finds more diverse and meaningful failure modes compared to baseline methods, and that fine-tuning VLAs on the generated instructions improves task success rates. Furthermore, results from a user study highlight that Q-DIG generates prompts judged to be more natural and human-like than those from baselines. Finally, real-world evaluations of Q-DIG prompts show results consistent with simulation, and fine-tuning VLAs on the generated prompts further success rates on unseen instructions. Together, these findings suggest that Q-DIG is a promising approach for identifying vulnerabilities and improving the robustness of VLA-based robots. Our anonymous project website is at qdigvla.github.io.
@@ I Introduction
While prior methods have been developed to induce failures in VLAs via adversarial instructions, it is still challenging to generate realistic instructions that elicit a controllably diverse range of failures. In particular, Embodied Red Teaming (ERT) [undefj] leverages in-context learning [undefk] to probe robot policies in simulation and generate adversarial instructions. Although ERT can uncover failure cases, it does not explicitly target a set of failure modes specified by system designers. Additionally, some generated instructions may fall outside the distribution of instructions that users normally provide. For the Coke can task above, one such unrealistic instruction would be “spot the red and white can lying flat and push it with precision.”
@@
Instead of generating maximally adversarial yet unrealistic instructions, one method to promote realism is to anchor instruction generation to a set of user-defined “attack styles.” For example, Rainbow Teaming [undefl] defined a range of attack styles such as slang and technical terms, and encouraged adversarial instructions to be generated following these style guidelines. To prevent instruction search from collapsing on a narrow range of styles, Rainbow Teaming used Quality Diversity (QD) optimization [undefm, undefn], which is a class of evolutionary algorithms that searches for high-quality solutions (in this case adversarial instructions) while maintaining diversity. However, although Rainbow Teaming has been effective at red-teaming LLMs, it solely operates in language space and does not consider visual context, which would be crucial for red-teaming VLAs.
@@
Inspired by these works, we present Q-DIG (Figure LABEL:fig:pull), a novel framework for red-teaming and improving VLAs. Our key insight is to use QD optimization to generate diverse adversarial instructions targeting the base VLA, while a VLM ensures that the generated instructions respect the visual context. Starting from an initial task instruction, Q-DIG iteratively mutates instructions, using previously discovered instructions as stepping stones, to build a repertoire of failure-inducing instructions across semantic categories, or “attack styles” [undefl]. After creating adversarial instructions, Q-DIG combines them with existing demonstrations to create an augmented dataset used to finetune the VLA, improving its robustness against diverse and adversarial instructions.
@@
Our contributions are: (1) We present Q-DIG, a framework that leverages Quality Diversity optimization for diverse and in-distribution adversarial instruction generation. (2) We evaluate Q-DIG in two simulation domains, SimplerEnv [undefo] and LIBERO [undefp], and show that compared to prior VLA red-teaming methods, Q-DIG produces more diverse and in-distribution adversarial instructions. (3) Through a user study, we show that instructions generated by Q-DIG are more human-like compared to those generated by prior methods. (4) We demonstrate that fine-tuning with the augmented dataset containing adversarial instructions improves robustness to unseen instructions. (5) We validate our approach in a sim-to-real setting, showing that the benefits of adversarial fine-tuning transfer to real-world robotic deployments.
@@ II-A Vision-Language-Action (VLA) Models
Vision-Language Models (VLMs) are a type of foundation model that integrates visual and language inputs to support multimodal reasoning. VLAs fine-tune large pre-trained VLMs to produce robot actions and thus inherit rich perceptual priors. Recent VLAs like RT-2 [undefq], OpenVLA [undefd], $\pi_{0}$ [undefe], and $\pi_{0.5}$ [undeff] highlight VLAs’ potential for generalization across robotic tasks. These models are trained on large-scale robotics datasets [undefr, undefs, undefc] and fine-tuned [undeft] to adapt to new domains or hardware.
Despite this progress, VLAs remain brittle in practice: small changes in task phrasing can induce failures, even when instructions are semantically equivalent [undefi]. In part, this arises because current VLA datasets and fine-tuning protocols provide only a narrow mapping between language instructions and demonstrations. For example, LIBERO [undefp] provides a single instruction per task with multiple demonstrations, DROID [undefs] associates each trajectory with at most three instructions, and Bridge [undefu] has a one-to-one mapping between instructions and demonstrations. We address this limitation by generating realistic adversarial instructions to expose failure modes in VLAs, followed by pairing them with demonstrations and retraining VLAs to mitigate them.
@@ II-B Red-Teaming Foundation Models
Foundation models can produce harmful behaviors, ranging from reinforcing societal biases and generating offensive content to leaking private information [undefh, undefv, undefw]. While filtering, fine-tuning, and reinforcement learning from human feedback [undefx, undefv] partially mitigate these issues, models remain vulnerable to targeted attacks. Such efforts are often driven by red-teaming (also called jailbreaking, if the intent is to cause harm), where adversarial inputs elicit unexpected outputs to uncover hidden vulnerabilities [undefh, undefy, undefz].
Robots controlled by foundation models are also prone to jailbreaking, with potential physical safety consequences. Prior work [undefg] showed that jailbroken LLM-controlled robots can be induced to take harmful actions across self-driving, unmanned vehicle navigation, and quadruped control. Prior research has also identified such vulnerabilities by predicting performance degradation under environmental variations [undefaa] and by searching over geometric object perturbations [undefab].
@@ II-C Quality Diversity for Foundation Models
Quality Diversity (QD) algorithms search for a diverse collection of high-performing solutions [undefac] by leveraging previously found solutions as stepping stones to generate new solutions that are higher performing or more diverse. QD has been applied in robotics, including training diverse RL agents [undefad, undefae, undefaf, undefag], and more recently extended to text generation with LLMs [undefah, undefai, undefaj, undefak], where previously discovered prompts [undefai] or poems [undefaj] serve as context for an LLM to generate new and improved variants.
The QD works closest to ours are Rainbow Teaming [undefl] and PLAN-QD [undefal]. Rainbow Teaming treats adversarial prompt generation as a QD problem, using LLMs as mutators that generate prompt variants and critics that evaluate outputs to discover diverse failures in language tasks. It also uses semantic failure modes, or “attack styles” (e.g., slang or technical terms) as measures to diversify prompts along user-relevant styles. Conditioning on these attack styles makes generated prompts more realistic. On the evaluation side, PLAN-QD searches for diverse prompts that induce varied collaborative behaviors in LLM-powered agents for sequential decision-making tasks. While both Rainbow Teaming and PLAN-QD advance systematic prompt search, they lack grounding in physical tasks and visual information required for embodied domains. In contrast, our method leverages visual grounding and feedback from rollouts of an embodied agent to generate diverse adversarial instructions.
@@ IV Method: Q-DIG
To generate diverse task instructions for VLAs, we propose Q-DIG (Quality Diversity for Diverse Instruction Generation), a framework that integrates Quality Diversity (QD) optimization with VLAs. The central idea is to use QD to automatically generate a wide range of semantically diverse instructions that, when executed on the base VLA, serve as adversarial prompts revealing failure modes. In this section, we first describe how we formulate generating adversarial instructions for VLAs as a QD problem, then provide the detailed framework of QD optimization, and finally explain how the resulting instructions can be leveraged for fine-tuning. Figure 1 presents an overview of our pipeline.
@@ IV-B Generating Adversarial Instructions
Instruction Mutation: Extending prior work in LLM jailbreaking [undefl], Q-DIG leverages a VLM as a mutator. Given an existing adversarial instruction $\bm{c}$}, its attack style $m({\bm{c}})$, initial observation for the task ${\bm{o}}_{0}$ (visual), and a target attack style category $z$, the mutator VLM leverages in-context learning [undefk] to generate a candidate instruction in the target attack style (e.g., “use of adverbs” in Fig. 1). We further generate $k$ sets of candidate instructions of batch size $b$ each (where $b$ and $k$ are hyperparameters). We then project each generated instruction into an embedding space with Sentence-BERT [undefam] and compute the average pairwise cosine similarity for each generated set. Among the generated sets, we select the one with the the most diverse set of instructions, i.e. the least pairwise similarity.
@@ IV-C Fine-Tuning the VLA
Following prior work [undeft], we perform supervised fine-tuning of the base VLA $\pi({\bm{o}},{\bm{c}})$ with data from ${\mathcal{D}}_{\mathrm{aug}}$, but the augmented dataset by itself is agnostic to the fine-tuning method. In summary, we use the diverse adversarial instructions generated by Q-DIG to enrich the training dataset. Once finetuned on this dataset, the VLA becomes more robust to different ways of describing the task.
@@ V-A Generating Adversarial Instructions
Domain: We evaluate Q-DIG in SimplerEnv [undefo] and LIBERO [undefp], two standard simulation domains for evaluating generalist robot policies. We use five tasks in SimplerEnv (picking up a Coke can, an apple, and a sponge, and opening and closing the top and bottom drawers) and all ten in the LIBERO-Goal task suite (different goals from a common starting setup, e.g., turning on the stove, putting a bowl on the stove, etc.) to highlight the effect of language instructions on VLA performance. We tested on other suites in LIBERO (e.g., LIBERO-Spatial, where objects were placed in different spatial arrangements for different tasks) but found that VLAs overfit to visual information of those tasks. Due to the one-to-one mapping between the starting observation and the task, VLAs solved those tasks with only image observations.
To ensure a good baseline performance of VLAs on these tasks before adversarial instruction generation, we fine-tuned them with the provided demonstration dataset for LIBERO and demonstrations collected via teleoperation for SimplerEnv. For both domains, we selected OpenVLA [undefd] and followed the process outlined by OpenVLA-OFT [undeft] for fine-tuning the base VLA for each domain. For LIBERO, we also conducted the same experiments on $\pi_{0.5}$ [undeff] and GR00T N1.6 [undefan]. We limited these additional VLAs to LIBERO due to time and computational constraints.
@@
Metrics: We compare the diversity of instructions generated by Q-DIG and the baselines with four diversity metrics: (1) BERT diversity, (2) BLEU diversity, (3) dissimilarity to the original instruction, and (4) archive coverage. The first two metrics, adopted from prior work [undefao, undefap, undefj], calculate the average pairwise similarity of the BERT embeddings [undefam] and the average pairwise BLEU score [undefaq], respectively. The third metric measures the average cosine similarity between the BERT embeddings of the original instructions and the generated instructions. We report $1-\text{similarity}$}” for these three diversity metrics. The fourth is a QD metric [undefac] that maps the generated instructions to the attack styles $\mathcal{Z}$} and computes the fraction of categories that have an instruction mapping to them (i.e., fraction of filled archive cells). Besides the diversity metrics, we compared the average failure variance induced by the generated instructions as a proxy for whether the instructions are within the VLA’s capability.
@@ VII Conclusion and Limitations
One limitation of Q-DIG is that it evaluates the success rate induced for a given instruction by rolling out the VLA in the environment multiple times. Hence, since VLA rollouts are computationally expensive, we were only able to run Q-DIG for 3–12 iterations. Another limitation is that the instructions are not generated with feedback from VLA training (i.e., leveraging fine-tuning losses), and hence, do not maximize learning potential. In the future, we will explore ways to reduce the computation required through approaches such as surrogate modeling [undefar], and thus enable even more scalable red-teaming of VLAs. Overall, we hope that our work paves the way for future work in scalable red-teaming methods for robust generalist robots.

### 03 Apr 2026 v3

@@ Abstract
Vision-Language-Action (VLA) models have significant potential to enable general-purpose robotic systems for a range of vision-language tasks. However, the performance of VLA-based robots is highly sensitive to the precise wording of language instructions, and it remains difficult to predict when such robots will fail. We propose Quality Diversity (QD) optimization as a natural framework for red-teaming embodied models, and present Q-DIG (Quality Diversity for Diverse Instruction Generation), which performs red-teaming by scalably identifying diverse, natural language task descriptions that induce failures while remaining task-relevant. Q-DIG integrates QD techniques with Vision-Language Models (VLMs) to generate a broad spectrum of adversarial instructions that expose meaningful vulnerabilities in VLA behavior. Our results across multiple simulation benchmarks show that Q-DIG finds more diverse and meaningful failure modes compared to baseline methods, and that fine-tuning VLAs on the generated instructions improves task success rates. Furthermore, results from a user study highlight that Q-DIG generates prompts judged to be more natural and human-like than those from baselines. Finally, real-world evaluations of Q-DIG prompts show results consistent with simulation, and fine-tuning VLAs on the generated prompts further success rates on unseen instructions. Together, these findings suggest that Q-DIG is a promising approach for identifying vulnerabilities and improving the robustness of VLA-based robots. Our anonymous project website is at qdigvla.github.io.
@@ I Introduction
We study this problem through the lens of red-teaming: the systematic discovery of inputs that expose failures in a model. While prior methods have been developed to induce failures in VLAs via adversarial instructions, it is still challenging to generate realistic instructions that elicit a controllably diverse range of failures. In particular, Embodied Red Teaming (ERT) [undefj] leverages in-context learning [undefk] to probe robot policies in simulation and generate adversarial instructions. Although ERT can uncover failure cases, it does not explicitly target a set of failure modes specified by system designers. Additionally, some generated instructions may fall outside the distribution of instructions that users normally provide. For the Coke can task above, one such unrealistic instruction would be “spot the red and white can lying flat and push it with precision.”
@@
We argue that Quality Diversity (QD) optimization [undefl, undefm] provides a natural framework for red teaming embodied models: it searches for high-performing solutions while maintaining diversity across user-defined dimensions. Prior work on algorithmic scenario generation has used QD to discover diverse failure modes in planning-based robotic systems [undefn, undefo]. Later work such as Rainbow Teaming [undefp] has applied QD to generate diverse adversarial prompts for large language models by targeting different “attack styles,” such as slang and technical terms, and encouraged adversarial instructions to be generated following these style guidelines. However, although Rainbow Teaming has been effective at red-teaming LLMs, it solely operates in language space and does not consider visual context, which would be crucial for red-teaming VLAs.
@@
Building on this formulation, we present Q-DIG (Figure LABEL:fig:pull), a novel framework for red-teaming and improving VLAs. Our key insight is that combining quality-diversity optimization with vision-language grounding enables the systematic discovery of diverse, realistic, and task-relevant failure modes in embodied systems. Starting from an initial task instruction, Q-DIG iteratively mutates instructions, using previously discovered instructions as stepping stones, to build a repertoire of failure-inducing instructions across semantic categories, or “attack styles” [undefp]. After creating adversarial instructions, Q-DIG combines them with existing demonstrations to create an augmented dataset used to finetune the VLA, improving its robustness against diverse and adversarial instructions.
@@
Our contributions are: (1) We present Q-DIG, a framework that leverages Quality Diversity optimization for diverse and in-distribution adversarial instruction generation. (2) We evaluate Q-DIG in two simulation domains, SimplerEnv [undefq] and LIBERO [undefr], and show that compared to prior VLA red-teaming methods, Q-DIG produces more diverse and in-distribution adversarial instructions. (3) Through a user study, we show that instructions generated by Q-DIG are more human-like compared to those generated by prior methods. (4) We demonstrate that fine-tuning with the augmented dataset containing adversarial instructions improves robustness to unseen instructions. (5) We validate our approach in a sim-to-real setting, showing that the benefits of adversarial fine-tuning transfer to real-world robotic deployments.
@@ II-A Vision-Language-Action (VLA) Models
Vision-Language Models (VLMs) are a type of foundation model that integrates visual and language inputs to support multimodal reasoning. VLAs fine-tune large pre-trained VLMs to produce robot actions and thus inherit rich perceptual priors. Recent VLAs like RT-2 [undefs], OpenVLA [undefd], $\pi_{0}$ [undefe], and $\pi_{0.5}$ [undeff] highlight VLAs’ potential for generalization across robotic tasks. These models are trained on large-scale robotics datasets [undeft, undefu, undefc] and fine-tuned [undefv] to adapt to new domains or hardware.
Despite this progress, VLAs remain brittle in practice: small changes in task phrasing can induce failures, even when instructions are semantically equivalent [undefi]. In part, this arises because current VLA datasets and fine-tuning protocols provide only a narrow mapping between language instructions and demonstrations. For example, LIBERO [undefr] provides a single instruction per task with multiple demonstrations, DROID [undefu] associates each trajectory with at most three instructions, and Bridge [undefw] has a one-to-one mapping between instructions and demonstrations. We address this limitation by generating realistic adversarial instructions to expose failure modes in VLAs, followed by pairing them with demonstrations and retraining VLAs to mitigate them.
@@ II-B Red-Teaming Foundation Models
Foundation models can produce harmful behaviors, ranging from reinforcing societal biases and generating offensive content to leaking private information [undefh, undefx, undefy]. While filtering, fine-tuning, and reinforcement learning from human feedback [undefz, undefx] partially mitigate these issues, models remain vulnerable to targeted attacks. Such efforts are often driven by red-teaming (also called jailbreaking, if the intent is to cause harm), where adversarial inputs elicit unexpected outputs to uncover hidden vulnerabilities [undefh, undefaa, undefab].
Robots controlled by foundation models are also prone to jailbreaking, with potential physical safety consequences. Prior work [undefg] showed that jailbroken LLM-controlled robots can be induced to take harmful actions across self-driving, unmanned vehicle navigation, and quadruped control. Prior research has also identified such vulnerabilities by predicting performance degradation under environmental variations [undefac] and by searching over geometric object perturbations [undefad].
@@ II-C Quality Diversity for Foundation Models
Quality Diversity (QD) algorithms search for a diverse collection of high-performing solutions [undefae] by leveraging previously found solutions as stepping stones to generate new solutions that are higher performing or more diverse. QD has been applied in robotics, including training diverse RL agents [undefaf, undefag, undefah, undefai]. QD has also been applied to generate diverse scenarios for autonomous agents [undefaj] and deployed robotic systems [undefn, undefo]. More recently, previous work has applied QD to text generation with LLMs [undefak, undefal, undefam, undefan], where previously discovered prompts [undefal] or poems [undefam] serve as context for an LLM to generate new and improved variants,
The QD works closest to ours are Rainbow Teaming [undefp] and PLAN-QD [undefao]. Rainbow Teaming extends scenario generation with QD to adversarial prompt generation, using LLMs as mutators that generate prompt variants and critics that evaluate outputs to discover diverse failures in language tasks. It also uses semantic failure modes, or “attack styles” (e.g., slang or technical terms) as measures to diversify prompts along user-relevant styles. Conditioning on these attack styles makes generated prompts more realistic. On the evaluation side, PLAN-QD searches for diverse prompts that induce varied collaborative behaviors in LLM-powered agents for sequential decision-making tasks. While both Rainbow Teaming and PLAN-QD advance systematic prompt search, they lack grounding in physical tasks and visual information required for embodied domains. In contrast, our method leverages visual grounding and feedback from rollouts of an embodied agent to generate diverse adversarial instructions.
@@ IV Method: Q-DIG
To generate diverse task instructions for VLAs, we propose Q-DIG (Quality Diversity for Diverse Instruction Generation), a framework that instantiates red-teaming of VLAs as a Quality Diversity (QD) optimization problem.
The central idea is to use QD to automatically generate a wide range of semantically diverse instructions that, when executed on the base VLA, serve as adversarial prompts revealing failure modes. In this section, we first describe how we formulate generating adversarial instructions for VLAs as a QD problem, then provide the detailed framework of QD optimization, and finally explain how the resulting instructions can be leveraged for fine-tuning. Figure 1 presents an overview of our pipeline.
@@ IV-B Generating Adversarial Instructions
Instruction Mutation: Extending prior work in LLM jailbreaking [undefp], Q-DIG leverages a VLM as a mutator. Given an existing adversarial instruction $\bm{c}$}, its attack style $m({\bm{c}})$, initial observation for the task ${\bm{o}}_{0}$ (visual), and a target attack style category $z$, the mutator VLM leverages in-context learning [undefk] to generate a candidate instruction in the target attack style (e.g., “use of adverbs” in Fig. 1). We further generate $k$ sets of candidate instructions of batch size $b$ each (where $b$ and $k$ are hyperparameters). We then project each generated instruction into an embedding space with Sentence-BERT [undefap] and compute the average pairwise cosine similarity for each generated set. Among the generated sets, we select the one with the the most diverse set of instructions, i.e. the least pairwise similarity.
@@ IV-C Fine-Tuning the VLA
Following prior work [undefv], we perform supervised fine-tuning of the base VLA $\pi({\bm{o}},{\bm{c}})$ with data from ${\mathcal{D}}_{\mathrm{aug}}$, but the augmented dataset by itself is agnostic to the fine-tuning method. In summary, we use the diverse adversarial instructions generated by Q-DIG to enrich the training dataset. Once finetuned on this dataset, the VLA becomes more robust to different ways of describing the task.
@@ V-A Generating Adversarial Instructions
Domain: We evaluate Q-DIG in SimplerEnv [undefq] and LIBERO [undefr], two standard simulation domains for evaluating generalist robot policies. We use five tasks in SimplerEnv (picking up a Coke can, an apple, and a sponge, and opening and closing the top and bottom drawers) and all ten in the LIBERO-Goal task suite (different goals from a common starting setup, e.g., turning on the stove, putting a bowl on the stove, etc.) to highlight the effect of language instructions on VLA performance. We tested on other suites in LIBERO (e.g., LIBERO-Spatial, where objects were placed in different spatial arrangements for different tasks) but found that VLAs overfit to visual information of those tasks. Due to the one-to-one mapping between the starting observation and the task, VLAs solved those tasks with only image observations.
To ensure a good baseline performance of VLAs on these tasks before adversarial instruction generation, we fine-tuned them with the provided demonstration dataset for LIBERO and demonstrations collected via teleoperation for SimplerEnv. For both domains, we selected OpenVLA [undefd] and followed the process outlined by OpenVLA-OFT [undefv] for fine-tuning the base VLA for each domain. For LIBERO, we also conducted the same experiments on $\pi_{0.5}$ [undeff] and GR00T N1.6 [undefaq]. We limited these additional VLAs to LIBERO due to time and computational constraints.
@@
Metrics: We compare the diversity of instructions generated by Q-DIG and the baselines with four diversity metrics: (1) BERT diversity, (2) BLEU diversity, (3) dissimilarity to the original instruction, and (4) archive coverage. The first two metrics, adopted from prior work [undefar, undefas, undefj], calculate the average pairwise similarity of the BERT embeddings [undefap] and the average pairwise BLEU score [undefat], respectively. The third metric measures the average cosine similarity between the BERT embeddings of the original instructions and the generated instructions. We report $1-\text{similarity}$}” for these three diversity metrics. The fourth is a QD metric [undefae] that maps the generated instructions to the attack styles $\mathcal{Z}$} and computes the fraction of categories that have an instruction mapping to them (i.e., fraction of filled archive cells). Besides the diversity metrics, we compared the average failure variance induced by the generated instructions as a proxy for whether the instructions are within the VLA’s capability.
@@ VII Conclusion and Limitations
One limitation of Q-DIG is that it evaluates the success rate induced for a given instruction by rolling out the VLA in the environment multiple times. Hence, since VLA rollouts are computationally expensive, we were only able to run Q-DIG for 3–12 iterations. Another limitation is that the instructions are not generated with feedback from VLA training (i.e., leveraging fine-tuning losses), and hence, do not maximize learning potential. In the future, we will explore ways to reduce the computation required through approaches such as surrogate modeling [undefo], and thus enable even more scalable red-teaming of VLAs. Overall, we hope that our work paves the way for future work in scalable red-teaming methods for robust generalist robots.